Opened 5 years ago

Closed 3 years ago

#558 closed flight processing (fixed)

MA14/21, flight day 296/2014, Sepilok

Reported by: dap Owned by:
Priority: whenever Milestone:
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by dap)

Data location: ~arsf/arsf_data/2014/flight_data/malaysia/MA14_21-2014_296_Sepilok

Data arrived from ARSF via network transfer on 05/01/2015.

Scientific objective: Quantifying the accuracy of current estimates of forest biomass by combining data retrieved from state-of-the-art ground-based laser scanner with airborne data.

PI: Mathias Disney

All Fenix header files have large differences between their fps_set and fps_qpf values (differences are between 0.278 and 0.673).

Basestation data are in DBX format.

OWL data were not collected.

Sensors:

  • Fenix (requested)
  • LiDAR (requested)
  • FW LiDAR (requested)
  • RCD (not requested)

Change History (31)

comment:1 Changed 5 years ago by dap

  • Description modified (diff)

comment:2 Changed 5 years ago by dap

PI would like data to be delivered by USB hard drive and FTP.

comment:3 Changed 4 years ago by dac

Basestation Processing

Basestation data looked like it contained two locations. Subset to GPS time 345600 - 354600 in GrafNav. Subset data processed OK. Antenna height wasn't set.

Lat 5 53 54.43758
Lon 118 03 55.38562
El. Height 65.953

comment:4 Changed 4 years ago by tec

Navigation Processing
IPAS TC still not happy, now complaining that sta.gpb does not exist.
Processed in grafnav, el mask set to 14.0deg.
Nav done. Copying back now

comment:5 Changed 4 years ago by tec

RCD Processing
Finished, ready for DC

comment:6 Changed 4 years ago by lah

RCD
Delivery check completed - ready for delivery

comment:7 Changed 4 years ago by dac

LiDAR Processing

Started lidar processing.

comment:8 Changed 4 years ago by dap

Starting Fenix Processing

comment:9 Changed 4 years ago by dap

  • Description modified (diff)

comment:10 Changed 4 years ago by dap

Removed first flightline from logsheet as it was aborted.

comment:11 Changed 4 years ago by stgo

Lidar Processing
Have taken over from dac.

Pitch and roll values seem consistent:

pitch -0.00625
roll -0.0024
heading 0.00008

Have had to recreate las files as alspp applied a 10 million metre offset. Will look in to the cause of this.

comment:12 Changed 4 years ago by asm

Fenix Processing
Re-started fenix processing.

comment:13 Changed 4 years ago by asm

Fenix Processing

There are 19 flightlines, flightline 20 does not have fenix data (as commented on the logsheet).

Most of the flightlines has a high number of clouds and present some issues (residuary bubbles, not overlapping features) on one side. Dale has pointed out that it could be because the flightlines are larger than planned and the plane continued flying collecting data. There are a lot of duplicated flightlines and we are putting aside and not processing the worst ones: 1, 4, 5, 6, 7, 10, 15. Even though, some of the flightlines that we are processing have still some minor issues that could not be solved with any sct value.

Last edited 4 years ago by asm (previous) (diff)

comment:14 Changed 4 years ago by asm

Fenix Processing

Found SCT values:

Flightline FENIX
2 0.78
3 -1.00
8 0.80
9 -1.10
11 -0.55
12 0.50
13 0.80
14 -0.20
16 1.00
17 -0.40
18 1.15
19 -0.01

Awaiting fenix calibration to continue.

Last edited 4 years ago by asm (previous) (diff)

comment:15 Changed 4 years ago by stgo

Lidar Processing

Full waveform complete, awaiting delivery check.

Pitch and roll values follow:

Pitch-0.0024
Roll-0.00625

comment:16 Changed 4 years ago by dac

RCD

Hardlinked RCD delivery to FTP (arsf17). To be included in lidar delivery email.

comment:17 Changed 4 years ago by dac

LiDAR DC

Starting delivery check

comment:18 Changed 4 years ago by dac

LiDAR DC

  • Empty table left in with height offsets - needs to be removed and text checked.
  • las1.0 folder should be las1.2 - Fixed
  • ASCII files had UNIX line endings - Fixed
  • Height looks offset compared to ASTER - also checked against SRTM and same problem. demcompare.py reports mean diff as 0.68 m so probably OK.
  • Not sure if it is worth delivering the 1st file - it contains hardly any data.
  • laszip returned error 'reading waveform descriptor cross-check' on the first file the rest were fine.
  • No waveforms in waveviewer for 1st file. The rest are fine.

comment:19 Changed 4 years ago by stgo

Lidar processing
I've fixed the readme and removed flightline 01. Also reproduced the dem to reflect the new data set, ready for re-DC.

comment:20 Changed 4 years ago by dac

LiDAR DC

Looks OK - ready to delivery

comment:21 Changed 4 years ago by gej

Lidar and RCD delivered via FTP

comment:22 Changed 4 years ago by stgo

Starting Fenix reprocessing

comment:23 Changed 4 years ago by stgo

Fenix reprocessing

Ready for delivery check.

This delivery needs to be thoroughly checked for:

  • correct lines included
  • correct linenames from logsheet
  • data quality comments
  • spectra shifts

comment:24 Changed 4 years ago by lah

Fenix DC - spectra
Checked spectra of all lines and no major problems. There is a persistent spike in the mask region for this flight, not unlike many other flights. The spectral shift is within 5 nm compared to previous data (2014 169).

comment:25 Changed 4 years ago by gej

Fenix DC

Flightline SCTs look okay, screenshots look okay.
Have moved logsheet into delivery,
Read me needs creating: Data Quality remarks:
There is some cloud cover on all included lines.
Not much overlap regions on included lines.

Screenshots need renaming according to proj_tidy

Will continue with DC once readme is generated.

comment:26 Changed 4 years ago by stgo

Fenix reprocessing
Readme generated and included in delivery, ready for re-DC

comment:27 Changed 4 years ago by gej

Fenix DC
Read-me comments are fine.
Screenshots have been renamed, everything looks good.

Lah has checked spectra

Mapped files zipped

Ready for delivery

Last edited 4 years ago by gej (previous) (diff)

comment:28 Changed 4 years ago by gej

Fenix delivered on 4th November 2015 via ftp

comment:29 Changed 4 years ago by lah

Archiving
Checked and transferred 30/11/15.

comment:30 Changed 3 years ago by dac

Archiving

Raw data available from NEODC - can't see processed (possibly hidden due to embargo).

comment:31 Changed 3 years ago by asm

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.